OpenStudy (anonymous):

The bottom of a box is a rectangle whose length is 6 cm longer than its width. How long must the box be in order for the area of the bottom to equal 453.25 cm2? 24.5 cm 18.5 cm 12.5 cm 30.5 cm

OpenStudy (anonymous):

so l=w+6 so then w=l-6 A=lw so substitute w A=l(l-6) A=l^2-6l you can then input the answers you have to find which one gets the area you desire

OpenStudy (anonymous):

ohhh so a is the answer?

OpenStudy (anonymous):

you can also input it into A=l(l-6) since it might be easier to do the math and yes

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problem